OpenStudy (anonymous):

SQRT (3Y+10)=SQRT (Y+11)-1

OpenStudy (anonymous):

Is this what it is? \[\sqrt{3y+10} = \sqrt{y+11}-1\]

OpenStudy (anonymous):

If so here you go:  Answer: no solution